# DetailGen3D: Generative 3D Geometry Enhancement via Data-Dependent Flow Ken Deng1,2,3  Yuanchen Guo2  Jingxiang Sun1  Zixin Zou2  Yangguang Li2  Xin Cai4 Yanpei Cao2  Yebin Liu1  Ding Liang1,2

### v1

@@ Abstract
Modern 3D generation methods can rapidly create shapes from sparse or single views, but their outputs often lack geometric detail due to computational constraints. We present DetailGen3D, a generative approach specifically designed to enhance these generated 3D shapes. Our key insight is to model the coarse-to-fine transformation directly through data-dependent flows in latent space, avoiding the computational overhead of large-scale 3D generative models. We introduce a token matching strategy that ensures accurate spatial correspondence during refinement, enabling local detail synthesis while preserving global structure. By carefully designing our training data to match the characteristics of synthesized coarse shapes, our method can effectively enhance shapes produced by various 3D generation and reconstruction approaches, from single-view to sparse multi-view inputs. Extensive experiments demonstrate that DetailGen3D achieves high-fidelity geometric detail synthesis while maintaining efficiency in training.
@@ 1 Introduction
Obtaining high-quality 3D geometry has been a long-standing research focus in the fields of computer vision and graphics. High-quality 3D models are not only valuable in the film industry, video games, and virtual reality, but also play a crucial role in the rapidly advancing field of embodied intelligence, contributing significantly to simulation environments. Early approaches for high-quality multi-view stereo reconstruction rely on dense multi-view inputs [16, 14, 1, 94, 26, 7]. Although recent methods [75, 78, 25, 6] based on Neural Radiance Field (NeRF) [48] and 3D Gaussian Splatting (3DGS) [30] have improved the end-to-end performance, reconstructing high-quality geometry remains a challenging problem.
@@
Recent methods for 3D generation from sparse or single views have evolved into two main paradigms: 1) optimization using 2D generative models and 2) direct training with 3D data. The first approach leverages pre-trained 2D vision models - notably, DreamFusion  [54] introduced Score Distillation Sampling (SDS) to align rendered views with text-conditioned distributions. Related methods  [79, 65] have extended this paradigm, though they consistently face geometric inconsistency issues, such as the “Janus problem” where details conflict across viewpoints. The second paradigm, training with 3D data, has seen multiple technical advances. Multi-view diffusion approaches  [40, 37, 36, 61, 33, 55, 44, 46, 60, 74] achieve view consistency but struggle with geometric coherence in fine details. Feed-forward Large Reconstruction Models  [24, 90, 32, 76, 106, 39, 81, 102] offer impressive speed but face resolution constraints. While recent 3D diffusion models  [99, 104, 103, 84, 56, 100] demonstrate quality improvements through direct 3D training, computational demands continue to limit their achievable resolution and detail fidelity.
@@
While recent advances in 3D generation have shown promising results, the generated shapes often lack fine geometric details. Simply scaling up these generative models demands prohibitive computational resources. Traditional geometry refinement approaches that rely on high-resolution dense multi-view images are ill-suited for enhancing such generated shapes, as these additional inputs are typically unavailable. Furthermore, existing optimization-based methods that utilize normal or shading information [83, 97, 82, 87, 50] require precise texture alignment and struggle to preserve global shape coherence. To address these challenges, we propose DetailGen3D, a generative approach to 3D geometry refinement that learns to synthesize plausible fine-scale details directly from a collection of high-quality 3D shapes. By capturing the underlying geometric patterns through data-driven priors, our method can enhance coarse shapes while maintaining structural consistency, even in the presence of noise or incomplete information.
@@
To tackle geometry refinement tasks with a generative model, a straightforward approach is to train a coarse geometry conditioned diffusion model, as seen in 2D image restoration methods [3, 96]. However, this method can be resource-intensive and may not be optimal for 3D tasks. Instead, inspired by Fischer et al. [12], who achieved effective image super-resolution without relying on large generative models, we propose a more training-efficient strategy. Rather than learning a complex mapping from noise to fine details, we focus on directly modeling the transformation between coarse and fine geometry using data-dependent rectified flow [41]. This approach utilizes optimal transport, which provides a direct and structured mapping between coarse and fine data points. By exploiting this coupling, we eliminate the need for the random coupling of noise and fine details, thus significantly reducing the training cost. This method is better suited for scaling up, offering a more efficient path to geometry refinement at larger scales.
Specifically, we introduce a training method called token matching. For the framework of our method, we follow 3DShape2Vecset [99]. In the 3D VAE encoding process, the query points obtained by downsampling are cross-attentioned with the global point cloud to get tokens. After denoising by DiT, the predicted tokens are sent to the 3D VAE decoder to obtain the implicit field as a 3D representation. In our method, after obtaining ground-truth (GT) geometry’s query points, we apply a nearest-neighbor algorithm to map the GT geometry’ query points across the entire coarse geometry point cloud. This approach ensures robust one-to-one token correspondence in space, maintains computational efficiency, and supports diverse geometric complexities.
@@
Effective geometry refinement requires carefully designed training data. While high-quality 3D shapes are available, obtaining matching coarse-fine pairs poses a significant challenge. Simply applying traditional mesh simplification or smoothing operations often produces artifacts that differ fundamentally from the characteristics of generated shapes. To address this, we leverage an LRM-based model that reconstructs 3D geometry from sparse-view renderings of high-quality fine models. This approach naturally introduces neural network-induced variations in geometric details, closely mimicking the characteristics of shapes produced by generative models. These generated shapes naturally serve as coarse inputs for training.
@@ 2 Related Works
Optimization based Model Reconstruction. Optimization based multi-view reconstruction methods experienced a rapid development with the emergence of NeRF [48]. However, methods like NeRF and NeuS [75] rely on lengthy optimization processes and dense multi-view inputs to achieve detailed geometry. Although subsequent approaches aimed at reducing the number of required views [77, 4, 28, 31, 42, 45, 57, 73, 68, 92], exploring reflective object reconstruction [43, 71, 18], improving reconstruction speed [49, 63, 64, 5, 17, 78], have been introduced, their geometric performance remains relying on dense multi view as input. Following the development of 3DGS [30], a representation that present faster rendering speed but worse geometry quality, a series of works focused on geometric optimization have emerged, including  [20, 25, 47, 98, 6], yet they have not resolved the issues of geometry quality and reconstruction speed.
@@
3D generation using 3D data training. Finetuning image diffusion models using Objaverse [9], Objaverse-XL [10] to generate reasonably consistent multi-views with 3D consistency, which could then be input into multi-view reconstruction models to obtain 3D geometry  [40, 37, 36, 61, 33, 55, 44, 46, 60, 74]. While methods combining these models with SDS optimization for 3D modeling are still relatively slow, their feedforward multi-view generation has significantly improved both generation quality and speed. However, optimization-based methods for obtaining 3D models still require several minutes to complete. Additionally, the generated results are still unsatisfactory because they are struggling to ensure 3D consistency in details. Some methods trying to improve multi-view consistency [70, 91, 80, 89], but geometry quality is even worse than previous methods since only 2D supervision is used. Using triplane as 3D representation, large-scale reconstruction model  [24, 32, 76, 90, 72, 66, 81, 106, 93] utilizing transformer architectures and triplane representations with fixed-view multi-view inputs, enabling reconstruction in just a few seconds. However, the extensive training time and GPU memory requirements of large reconstruction models severely limit their resolution. Generative models using triplane as 3D representation [19, 101, 23, 62, 85] meet the same problem as well. Using point cloud as 3D representation improves the efficiency [29, 51, 99, 104, 103] also brings high training costs. Some recent methods use voxel [56, 38] or gaussian [70, 102, 100] as 3D representation are also efficient.
@@ 3 Methods
Figure 2: (1) Inference pipeline. We use FPS-VAE to extract tokens of the coarse geometry generated or reconstructed, then input the coarse token and DINO feature [52] of the image prompt to DiT [53]. After denoising, we decode the predicted token using FPS-VAE decoder to obtain refined geometry. The inference process takes only a few seconds. (2) For training data, we use reconstruction results reconstructed by LRM using multi-views rendered from fine geometry as coarse geometry. (3) We demonstrate the token matching process on the left. On the right, for the top one, we only use part query points, which are located in quadrant one, and for the bottom one, we use full query points, which demonstrate tokens represent the space around the corresponding query points.
We formulate our geometry enhancement task as follows: given an input of coarse geometry obtained by reconstruction or generation and a corresponding view of the desired fine geometry, the objective is to produce fine geometry that aligns with both the image prompt and the input coarse geometry. The inference pipeline is shown in Fig 2.
3.1 Model Architecture
Our network structure mainly consists of two parts: FPS-VAE and Image conditioned DiT. FPS-VAE encoder is used to encode 3D shapes into tokens to save GPU memory. Different from  [103], which directly downsample full point cloud to obtain query points, we add farthest point sampling (FPS) [15] process after random downsampling of full point cloud and the decoder is used to decode tokens into the implicit field. The Image conditioned DiT is used to model the refinement process, using coarse geometry’s token as input and predicting fine geometry’s token.
FPS-VAE. The design of our FPS-VAE is primarily inspired by 3DShape2VecSet [99] and CLAY [103]. To improve detail capture, we increase the point sampling count from CLAY’s maximum of 8192 to 20480. Additionally, we introduce a two-step downsampling process for the point cloud: the first step uses random downsampling, followed by FPS [15] downsampling. This approach helps ensure that different tokens correspond to distinct spatial regions, thereby improving the model’s ability to capture fine-grained details.
Image conditioned DiT. Our DiT network is designed with efficiency in mind, comprising 24 transformer layers with a width of 768, 12 attention heads, and a latent length of 2048, totaling 368M parameters. To accommodate the relatively small width, we inject the timestep using the adaLN [27] method:
$z=z+g_{\text{msa}}\cdot\text{self\_attn}(\text{mod}(\text{LN}_{\text{msa}}(z),s_{\text{msa}},\gamma_{\text{msa}}))$ (1)
Turning to conditioning on the image prompt, we use cross-attention, ensuring that the image guides the refinement process effectively while maintaining robust feature capture. The conditioning is defined as:
$z=z+g_{\text{mca}}\cdot\text{cross\_attn}(\text{mod}(\text{LN}_{\text{mca}}(z),s_{\text{mca}},\gamma_{\text{mca}}),y)$ (2)
@@ 3.1 Model Architecture
For encoding, we sample $N=20480$ surface points from a 3D mesh, which are then transformed into latent tokens $z$ using $\mathcal{E}(\cdot)$. Simultaneously, a pretrained ViT (DINO V2 [52]) processes the image prompt into a feature map $y$. The DiT then predicts the tokens for fine geometry, conditioned on these inputs, ensuring fine-grained detail capture during the refinement stage.
3.2 Data Dependent Rectified Flow
We use data-dependent rectified flow to model the distribution mapping of coarse and fine geometry due to its efficiency. The noise added to fine token is sampling from coarse distribution in the training stage and in the inference stage, the denoising process is begin with the sampling from coarse distribution. This modeling can natually almost predict coarse geometry without training. In contrast to cross attention which modeling the distribution mapping between noise and fine geometry, using coarse geometry as condition needs long time training to achieve noise to coarse mapping.
Rectified Flow and Loss Functions. Unlike previous generative models that model the mapping between a Gaussian distribution and the GT data distribution, we model the mapping between the coarse geometry distribution and the fine geometry distribution. Let $x_{1}$ represent the fine geometry and $x_{0}$ represent the coarse geometry. The geometry at time $t$, $x_{t}$, is given by:
$x_{t}=(1-t)\cdot x_{0}+t\cdot x_{1}$ (4)
The conditional probability $P(x_{1}|x_{t})$ is defined as:
$P(x_{1}|x_{t})=\mathcal{N}(x|(1-t)\cdot x_{0}+t\cdot x_{1},\sigma_{\min}^{2}\cdot I)$ (5)
@@ 3.2 Data Dependent Rectified Flow
$L(\theta)=\mathbb{E}_{t,x_{0},x_{1},y}\left[\left\|v_{\theta}(t,x_{t},y)-(x_{1}-x_{0})\right\|^{2}\right]$ (6)
where $v_{\theta}$ is parameterized by the DiT network discussed above with learnable parameters $\theta$, and $y$ is the image condition guiding the geometry refinement.
Noise Augmentation. The degraded geometric distribution is relatively independent so direct modeling of distribution transformation between coarse and fine geometry cannot well learn comprehensive refinement rules from the dataset, leading to noisy results. Inspired by [13], we apply noise augmentation to the token of coarse geometry, which helps improve the quality. During training, we apply noise to $x_{0}$ according to the DDPM [22] linear schedule at 400 timesteps. The noise $\epsilon$ is added to $x_{0}$ before obtaining $x_{t}$, following the DDPM [22] linear schedule. This is done because applying noise after obtaining $x_{t}$ would affect the efficiency of optimal transport, leading to an optimization path that is not straight. During inference stage, we also noise $x_{0}$ at 100 timesteps.

### v3

@@ Abstract
Modern 3D generation methods can rapidly create shapes from sparse or single views, but their outputs often lack geometric detail due to computational constraints. We present DetailGen3D, a generative approach specifically designed to enhance these generated 3D shapes. Our key insight is to model the coarse-to-fine transformation directly through data-dependent flows in latent space, avoiding the computational overhead of large-scale 3D generative models. We introduce a token matching strategy that ensures accurate spatial correspondence during refinement, enabling local detail synthesis while preserving global structure. By carefully designing our training data to match the characteristics of synthesized coarse shapes, our method can effectively enhance shapes produced by various 3D generation and reconstruction approaches, from single-view to sparse multi-view inputs. Extensive experiments demonstrate that DetailGen3D achieves high-fidelity geometric detail synthesis while maintaining efficiency in training. Our project page is https://detailgen3d.github.io/DetailGen3D/
@@ 1 Introduction
Obtaining high-quality 3D geometry has been a long-standing research focus in the fields of computer vision and graphics. High-quality 3D models are not only valuable in the film industry, video games, and virtual reality but also play a crucial role in the rapidly advancing field of embodied intelligence, contributing significantly to simulation environments. Early approaches for high-quality multi-view stereo reconstruction rely on dense multi-view inputs [19, 17, 1, 85, 27, 8]. Although recent methods [67, 69, 26, 5] based on Neural Radiance Field (NeRF) [44] and 3D Gaussian Splatting (3DGS) [30] have improved the end-to-end performance, reconstructing high-quality geometry remains a challenging problem.
@@
Recent methods for 3D generation from sparse or single views have evolved into two main paradigms: 1) optimization using 2D generative models and 2) direct training with 3D data. The first approach leverages pre-trained 2D vision models - notably, DreamFusion  [50] introduced Score Distillation Sampling (SDS) to align rendered views with text-conditioned distributions. Related methods  [70, 59] have extended this paradigm, though they consistently face geometric inconsistency issues, such as the “Janus problem” where details conflict across viewpoints. The second paradigm, training with 3D data, has seen multiple technical advances. Multi-view diffusion approaches  [40, 37, 36, 57, 32, 51, 42, 43, 56, 66] achieve view consistency but struggle with geometric coherence in fine details. Feed-forward Large Reconstruction Models  [25, 82, 31, 68, 99, 39, 72, 94] offer impressive speed but face resolution constraints. While recent 3D diffusion models  [91, 96, 95, 75, 53, 92] demonstrate quality improvements through direct 3D training, computational demands continue to limit their achievable resolution and detail.
@@
While recent advances in 3D generation have shown promising results, the generated shapes often lack fine geometric details. Simply scaling up these generative models demands prohibitive computational resources. Traditional geometry refinement approaches that rely on high-resolution dense multi-view images are ill-suited for enhancing such generated shapes, as these additional inputs are typically unavailable. Furthermore, existing optimization-based methods that utilize normal or shading information [74, 88, 73, 79, 45] require precise texture alignment and struggle to preserve global shape coherence. To address these challenges, we propose DetailGen3D, a generative approach to 3D geometry refinement that learns to synthesize plausible fine-scale details directly from high-quality 3D shapes. By capturing the underlying geometric patterns through data-driven priors, our method can enhance coarse shapes while maintaining structural consistency, even under noisy or incomplete information.
@@
To tackle geometry refinement tasks with a generative model, a straightforward approach is to train a coarse geometry conditioned diffusion model, as seen in 2D image restoration methods [3, 87]. However, this method can be resource-intensive and may not be optimal for 3D tasks. Instead, inspired by Fischer et al. [14], who achieved effective image super-resolution without relying on large generative models, we propose a more training-efficient strategy. Rather than learning a complex mapping from noise to fine details, we focus on directly modeling the transformation between coarse and fine geometry using data-dependent rectified flow [41]. This approach utilizes optimal transport, which provides a direct and structured mapping between coarse and fine geometry. By exploiting this coupling, we eliminate the need for the random coupling of noise and fine details, thus significantly reducing the training cost. This method is better suited for scaling up, offering a more efficient path to geometry refinement.
Specifically, we introduce a training method called token matching. Establishing the refinement process locally is essential; otherwise, global refinement leads to inefficient training and hinders detail capturing. Token matching matches the coarse geometry latent code with the fine geometry latent code in latent space one-to-one, improving training efficiency. It prevents the network from learning unnecessary operations (e.g., swapping positions between two latent codes) and focuses solely on local refinement operations. This approach enables the capture of geometry details even with a small network.
@@
Effective geometry refinement requires carefully designed training data. While high-quality 3D shapes are available, obtaining matching coarse-fine pairs poses a significant challenge. Simply applying traditional mesh degradation algorithms (e.g., simplification or smoothing operations) leads to simple objects remaining unchanged or extreme degradation on complex objects, resulting in low-quality coarse-fine pairs. To address this, we leverage an LRM-based model that reconstructs 3D geometry from sparse-view renderings of high-quality fine models. This approach enables consistent degradation across objects of varying complexity, enhances the utilization of existing 3D objects, and increases both the quality and quantity of coarse-fine pairs.
@@ 2 Related Works
3D generation using 3D data training. Finetuning image diffusion models using Objaverse [11], Objaverse-XL [12] to generate reasonably consistent multi-views with 3D consistency, which can then be input into multi-view reconstruction models to obtain 3D geometry  [40, 37, 36, 57, 32, 51, 42, 43, 56, 66]. Some methods attempting to improve multi-view consistency [63, 83, 71, 81], but geometry quality is even worse than previous methods since only 2D supervision is used. Using triplane as 3D representation, large-scale reconstruction model  [25, 31, 68, 82, 65, 60, 72, 99, 84] utilizing transformer architectures and triplane representations with fixed-view multi-view inputs, enabling reconstruction in just a few seconds. However, the extensive training time and GPU memory requirements of large reconstruction models severely limit their resolution. Generative models using triplane as 3D representation [21, 93, 24, 58, 76] meet the same problem as well. Using point cloud as 3D representation improves the efficiency [29, 46, 91, 96, 95, 97, 33] but also brings high training costs. Some recent methods use voxel [53, 38, 78] or gaussian [63, 94, 92] as 3D representations are also efficient.
3D shape detailization. SketchPatch [16] uses a PatchGAN [28] discriminator to mimic the local style of a reference image, stylizing plain solid-lined sketches. DECORGAN [10], which employs PatchGAN to generate detailed voxel shapes from input coarse voxels, with the geometric style of the generated shape derived from a detailed reference voxel model. ShaDDR [6] enhances the generated geometry of DECORGAN by utilizing a 2-level hierarchical GAN and introducing texture generation. Meanwhile, DECOLLAGE [7] further improves shape detailization through structure-preserving losses and adaptive weighting of style and global discriminators. However, these methods primarily focus on enhancing the voxel resolution of low-quality inputs, which is different from our method: pushing the boundaries of detail enhancement beyond current state-of-the-art generative models, capturing finer details missed by coarser methods.
@@ 3 Methods
Figure 2: (1) Inference pipeline. We use 3D-VAE to extract tokens of the coarse geometry generated or reconstructed, then input the coarse token and DINOv2 feature [47] of the image prompt to DiT [48]. After the refinement process, we decode the predicted token using a 3D-VAE decoder to obtain refined geometry. The inference process takes only a few seconds. (2) For training data, we use reconstruction results reconstructed by LRM using multi-views rendered from fine geometry as coarse geometry. (3) We demonstrate the token matching process on the left. On the right, for the top one, we only use part query points, which are located in quadrant one, and for the bottom one, we use full query points, which demonstrate that tokens represent the space around the corresponding query points.
Our method $F$ is designed to refine coarse geometry $G_{\text{Coarse}}$ into fine geometry $G_{\text{Fine}}$, guided by an image prompt $I$. Here, $G_{\text{Coarse}}$ may originate from reconstruction or generation processes, while $G_{\text{Fine}}$ represents an enhanced version with improved surface quality and additional geometric details, such as noise removal and refinement:
$G_{Fine}=\mathcal{F}(G_{Coarse},I)$ (1)
Considering this process has strong uncertainty, we model this process using a generative model. In particular, we select rectified flow [41] because of its efficiency based on optimal transport and its ability to model the mapping relationship between two data distributions, i.e., between the distribution of coarse geometry and fine geometry. Our inference pipeline is shown in Fig 2 (1).
3.1 Data Dependent Rectified Flow
We use data-dependent rectified flow to model the distribution mapping of coarse and fine geometry due to its efficiency. Compared to modeling the transformation between Gaussian noise and fine geometry, directly modeling the transformation between coarse and fine geometry can, to some extent, reconstruct the coarse shape without extensive denoising training, as demonstrated in Fig 7. In contrast, modeling the distribution mapping between noise and fine geometry, with coarse geometry as a condition via cross-attention, requires longer training time to achieve noise-to-coarse mapping.
Model Architecture. To improve efficiency, we build our refinement process in latent space, and our network consists of two parts: 3D Variational Autoencoder (3D-VAE) and Diffusion Transformer [48] (DiT).
3D Variational Autoencoder. The design of our 3D-VAE is primarily inspired by 3DShape2VecSet [91], CLAY [95], and TripoSG [33] and shares the same architecture with CLAY’s VAE. To transform 3D geometry into latent space, we first sample a point cloud $X$ from 3D geometry’s surface and adopt a two-stage downsampling process (random downsample firstly then apply farthest point sampling [18]) for $X$ to get query points $X_{0}$. Lastly, the cross attention is applied to $X$ and $X_{0}$ with learnable positional embeddings to obtain the latent code $z$.:
$z=\mathcal{E}(X)=\text{CrossAttn}(\text{PosEmb}(X_{0}),\text{PosEmb}(X))$ (2)
The 3D-VAE decoder, composed of several self attention layers and a cross attention layer, transforms the latent code $z$ to a signed distance field (SDF) with preset query points $q$.:
$SDF=\mathcal{D}(q,z)=\text{CrossAttn}(\text{PosEmb}(q),\text{SelfAttn}(z))$ (3)
Diffusion Transformer. Our DiT network, comprised of 24 DiT blocks with a width of 768 and totaling 368M parameters, is designed with efficiency in mind. Each DiT block consists of a multi-head cross-attention layer (MCA), a multi-head self-attention layer (MSA), and a feedforward network (FFN), interleaved with layer normalization (LN). To accommodate the relatively small width, we inject the time step $t$ using the adaptive layer normalization (adaLN) [49], modulating MSA, MCA, and FFN via factors $g$, $\gamma$, and $s$, obtained by a MLP conditioned on $t$.:
$z=z+g_{\text{msa}}\cdot\text{SelfAttn}(\text{mod}(\text{LN}_{\text{msa}}(z),s_{\text{msa}},\gamma_{\text{msa}}))$ (4)
Turning to conditioning on the image prompt $y$, we adopt cross-attention to ensure spatial alignment between latent code and image feature, which is extracted by DINOv2 [47]. The conditioning is defined as:
$z=z+g_{\text{mca}}\cdot\text{CrossAttn}(\text{mod}(\text{LN}_{\text{mca}}(z),s_{\text{mca}},\gamma_{\text{mca}}),y)$ (5)
@@ 3.1 Data Dependent Rectified Flow
Rectified Flow and Loss Functions. Unlike previous generative models modeling the mapping between a Gaussian distribution and the ground truth data distribution, we model the mapping between the coarse geometry distribution and the fine geometry distribution. Let $z_{1}$ represent the fine geometry’s latent code and $z_{0}$ corresponds to the coarse geometry’s latent code. The geometry latent code at time $t$, $z_{t}$, is given by:
$z_{t}=(1-t)\cdot z_{0}+t\cdot z_{1}$ (7)
The conditional probability $P(z_{1}|z_{t})$ is defined as:
$P(z_{1}|z_{t})=\mathcal{N}(z|(1-t)\cdot z_{0}+t\cdot z_{1},\sigma_{\min}^{2}\cdot I)$ (8)
@@
$L(\theta)=\mathbb{E}_{t,z_{0},z_{1},y}\left[\left\|v_{\theta}(t,z_{t},y)-(z_{1}-z_{0})\right\|^{2}\right]$ (9)
where $v_{\theta}$ is parameterized by the DiT with learnable parameters $\theta$ and $y$ is the image prompt. The L2 loss constrains the prediction of DiT’s prediction $v_{\theta}$ with respect to the ground truth velocity $(z_{1}-z_{0})$ in latent space.
Noise Augmentation. Since the degraded geometric distribution is relatively independent, direct modeling of distribution transformation between coarse and fine geometry cannot effectively learn comprehensive refinement rules from the dataset, leading to noisy results. Inspired by [15], we apply noise augmentation to the latent code of coarse geometry, which helps improve the quality. The noise $\epsilon$ is added to $z_{0}$ before obtaining $z_{t}$ instead of after, as applying noise after obtaining $z_{t}$ would affect the efficiency of optimal transport, leading to an optimization path that is not straight. See supplementary material for more details.